# RecJPQ: Training Large-Catalogue Sequential Recommenders

### 11 Dec 2023 v1

@@ Abstract.
Sequential recommender systems rank items based on the likelihood of their next appearance in user-item interactions. Current models such as BERT4Rec and SASRec generate sequence embeddings and compute scores for catalogue items, but the increasing catalogue size makes training these models costly. The Joint Product Quantisation (JPQ) method, originally proposed for passage retrieval, markedly reduces the size of the retrieval index with minimal effect on model effectiveness by replacing passage embeddings with a limited number of shared centroid embeddings. This paper introduces RecJPQ, a novel adaptation of JPQ for sequential recommendations. We apply RecJPQ to SASRec, BERT4Rec, and GRU4rec models on three large-scale sequential datasets. Our results showed that RecJPQ can notably reduce model size (e.g., 48$\times$ reduction for the Gowalla dataset with no effectiveness degradation). RecJPQ can also improve model performance through a regularisation effect (e.g. +0.96% NDCG@10 improvement on the Booking.com dataset).
journalyear: 2024copyright: acmcopyrightconference: 32nd ACM International Conference on Information and Knowledge Management; March 4–8, 2024; Mérida, Mexicobooktitle: WSDM ’24), March 4–8, 2024, Mérida, Mexico
@@ 1. Introduction
When a recommender has many items in the catalogue, various challenges arise in training the neural recommendation model. Firstly, the item embedding tensor may contain more model parameters than the rest of the model. For example, there are more than 800 million videos on YouTube [^2] https://earthweb.com/how-many-videos-are-on-youtube/. If the model uses 128-dimensional embeddings, the whole item embeddings tensor will have more than 100 billion parameters, which is comparable with the number of parameters of the largest available machine learning models (Brown et al., 2020), even without accounting for the parameters of the model’s intermediate layers. This is a problem specific to recommender systems. However, in the related area of dense passage retrieval (Karpukhin et al., 2020; Khattab and Zaharia, 2020), passage embeddings are obtained by encoding passage text using a pre-trained language model. In contrast, item side information, such as text, is not necessarily available in a typical recommender systems scenario; therefore, item embeddings should be directly learned from the interactions. Secondly, a large number of such trainable parameters also makes the model prone to overfitting. A third challenge caused by the large catalogue is the size of the output scores tensor (rightmost tensor in Figure 1): for example, in BERT4Rec, it contains a score for each item for every position, for every sequence in the training batch, so training BERT4Rec with more than 1 million items in the catalogue may be prohibitively expensive (Petrov and Macdonald, 2022a). This problem is typically solved using negative sampling, whereby instead of computing the full output tensor, the model computes scores for a small proportion of negative items not interacted by the user (e.g. SASRec (Kang and McAuley, 2018) uses one negative per positive). However, negative sampling comes with its own challenges (for example, it usually requires informative negative mining (Rendle, 2022)). Nevertheless, negative sampling is an orthogonal research direction, and in this paper, we use SASRec in cases where negative sampling is necessary. We refer to our recent publication (Petrov and Macdonald, 2023), where we analysed negative sampling problem in details. To summarise the challenges, a large embedding tensor increases the model size, slows model training down, and can necessitate further modelling tricks such as negative sampling, which bring their own challenges.
@@
There are some existing methods (Xia et al., 2023; Wang et al., 2022; Kang et al., 2020) for item embedding compression (we discuss these methods in Section 2). However, most of these methods compress the embedding tensor after the model is fully trained (including training the full embedding tensor). However, as argued above, training may be prohibitively expensive in large-scale recommender systems. Hence, this paper addresses the problem of a large item embedding tensor in sequential recommendation models at the training stage.
333
@@
To mitigate this problem, we propose a novel RecJPQ technique inspired by the success of a recent Joint Product Quantisation (JPQ) work (Zhan et al., 2021a) for text retrieval. JPQ itself is based on Product Quantisation (PQ) (Jégou et al., 2011), a popular method of compressing vectors by splitting them into sub-embeddings and encoding them using a discrete centroids codebook (the codebook maps from item ids to the associated centroids; see Section 3.1 for the details). The main innovation of JPQ compared to the standard PQ method is that it learns the centroids embeddings as part of the overall model training process. In contrast, PQ requires training the model first and only then compressing the embeddings (frequently, this second step uses external tools, such as FAISS (Johnson et al., 2021)). This means that JPQ does not need to keep the embedding matrix in memory during model training. We argue that this innovation is valuable for recommender systems. Indeed, as mentioned above, real-life recommender systems can have hundreds of millions of items in their catalogues and keeping full embeddings tensor in memory may be prohibitively expensive. This is particularly important for deep-learning-based sequential recommender systems because these models require keeping the whole model in GPU (or TPU) memory during training. GPU memory is costly even when compared to regular computer RAM.
@@
Unfortunately, it is hard to adapt JPQ to the recommendation scenario, as it is specific to textual information retrieval. In particular, JPQ assumes the existence of a pre-trained passage retrieval model and index, which it uses to assign items to centroids (see more details in Section 3.2). These pre-trained models rarely exist in item recommendations. Hence, in RecJPQ, we experiment with performing the initial assignment of centroids using three different strategies: (i) discrete truncated SVD (centroids obtained by discretising the item representations obtained by an SVD decomposition of the user-item matrix), (ii) discrete BPR (centroids obtained by discretising the item embeddings obtained from BPR) and (iii) random assignments. We describe these assignment strategies in detail in Section 4.
@@
RecJPQ is a model component that replaces traditional item embeddings in sequential recommender systems. In general, it can be applied to any recommender system based on item embeddings, but in this paper, we focus specifically on sequential models, as in these models, item embeddings comprise the biggest part of the model (e.g. sequential models usually do not have user embeddings). In contrast with existing methods, RecJPQ does not require training full uncompressed embedding and does not modify the original model loss function. Our experimentation on three datasets (see Section 5) demonstrates that RecJPQ can be successfully applied to different models, including SASRec (Kang and McAuley, 2018), BERT4Rec (Sun et al., 2019) and GRU (Hidasi et al., 2016; Petrov and Macdonald, 2022a), achieving a large factor of embeddings compression (e.g. 47.94x compression of SASRec on Gowalla) without any effectiveness degradation. Moreover, on 2 out of our 3 experimental datasets, applying RecJPQ increases model performance (e.g. +0.96% NDCG@10 on Booking.com dataset, significant improvement); we attribute these improvements to model regularisation.
In short, the contributions of this paper are as follows:
  1. (1)
    We propose RecJPQ, a novel technique for reducing the size of sequential recommendation models during training based on Joint Product Quantisation.
  2. (2)
    We propose three strategies for assigning codes to items, two of which (discrete truncated SVD and discrete BPR) assign similar codes to similar items, and one assigns codes randomly.
  3. (3)
    We perform an extensive experimental evaluation of RecJPQ on three datasets and show that RecJPQ allows reducing the models’ size without hindering the model performance.
The rest of the paper is organised as follows: Section 2 introduces related work on embeddings compression and identifies limitations of existing methods; Section 3 covers Product Quantisation (PQ) and Joint Product Quantisation (JPQ) - the methods, which serve as the basis for our work; Section 4 introduces RecJPQ and covers centroid assignment strategies for RecJPQ; in Section 5 we experimentally evaluate RecJPQ; Section 6 contains final remarks.
@@ 2. Related work
Table 1. Existing embedding compression methods. Desired method characteristics are highlighted in bold.
@@
This section covers existing work on compressing and discretising embeddings in recommender systems, identifies the limitations in existing work and positions our contributions in the context of existing methods. Table 1 summarises existing methods and positions RecJPQ, our proposed compression technique. The table highlights with boldface the desirable characteristics necessary for training a large-scale [^5] For simplicity, we say that a catalogue is ”large-scale” if it has more than 1 million, as it becomes challenging to train recommender systems on that scale (Petrov and Macdonald, 2022a). sequential model, specifically: the method can be applied to work with different backbone sequential models, and does not require training full embeddings (as we work with the assumption that full embeddings tensor does not fit into GPU memory); and we want the model to focus on item embeddings rather than embeddings of other entities, such as users or features. As illustrated in the table, the methods for compressing the models can be broadly divided into two groups: model dependent and model agnostic.
666
In the model-dependent methods (Lian et al., 2020; Xia et al., 2023), embedding compression mechanism is integrated as a component into the recommendation model itself. Hence the training architecture of these methods has to be aware of the compression, and the loss function includes components responsible for the embedding compression. For example, LightRec (Lian et al., 2020) uses the Deep Semantic Similarity Model (DSSM) (Huang et al., 2013) as the backbone model and uses an additional knowledge distillation component in the loss, which allows for learning compressed representations of the embeddings. However, DSSM is not a sequential model, and it is unclear whether or not the method can be adapted to the sequential recommendation case. Similarly, EODRec (Xia et al., 2023) which uses SASRec (Kang and McAuley, 2018) as its backbone, one of the most popular sequential models based on the Transformer architecture (Vaswani et al., 2017). The loss function of EODRec also consists of four components, some of which are responsible for recommendation, and others are responsible for embedding compression. While SASRec, used by EODRec as its backbone, is an efficient model, in many cases other models such as BERT4Rec show better results (Sun et al., 2019; Petrov and Macdonald, 2022b). In general, while some model-dependent methods may substantially reduce the size of a trained model, these methods have several limitations, which make them unsuitable for training sequential recommendation models with large catalogues. In particular:
@@
L2 Model-dependent methods usually require training (uncompressed) item embeddings and then use knowledge distillation or teacher-student techniques to obtain compressed representations of the embeddings. This approach substantially reduces the final model size, thereby helping inference on smaller devices, but requires a large amount of GPU memory while training, thereby limiting the overall number of items in the catalogue. For this reason, the main positioning for EODRec model (Xia et al., 2023) is the on-device recommendation: while the final model produced by this method is small, it requires storing full item embeddings while training. Post-training quantisation methods (Yao et al., 2022), which recently became popular to reduce the size of large language models via quantising their weights into lower-precision numbers (e.g. float16, or int8) also have this limitation – they need to have access to full model before quantising. Similarly, Mixed Precision Training (Micikevicius et al., 2018) builds a smaller precision model, but it requires keeping full precision weights in memory. Placing the embeddings tensor into Approximate Nearest Neighbours (Johnson et al., 2021) or Hierarchical Navigable Small Worlds (Malkov and Yashunin, 2020) indexes also requires access to the full embedding tensor at model training time, and therefore also exhibits this limitation.
@@
On the other hand, the existing model-agnostic methods (Shi et al., 2020a, b) do not depend on the specific model architecture, and likewise do not add extra components to the loss functions. Typically, these methods implement a mechanism that takes the place of the embeddings tensor in the backbone model, and hence can be used with many models. However, on inspection of the relevant work, we identified additional limitations of these methods:
777
@@
L4 Many methods are not designed for compressing item embeddings. For example, PreHash (Shi et al., 2020a) is a method specific for compressing user embeddings (i.e. it uses the user’s history to construct user embeddings). The method uses an attention network over the history of user interactions. Adapting this network structure for items embeddings is a hard task: a user may only interact with a few items; in contrast, a popular item may be interact with by millions of users. The attention mechanism depends quadratically on the sequence length, and therefore applying it to users who interacted with a popular item would be prohibitively expensive.
@@
L5 Finally, many methods lack structure in compressed item representations. This leads to situations where unrelated items have similar representations and, conversely, when similar items have dissimilar representations. Both these cases may limit the models’ generalisation ability and hinder the models’ performance. One example of such unstructured methods is hashing-based Quotient Remainder (Shi et al., 2020b), which compresses embeddings of categorical features (e.g., genres). Quotient Remainder assigns feature codes based on the quotient and the remainder of the division of the feature id by some number. When applied to item ids (items can also be seen as categorical features), the quotient and the remainder are unrelated to the item characteristics. Hence similar items are unlikely to have similar codes. Nevertheless, Quotient Remainder is one of the few methods that can be used to train a model on a large-scale dataset, and therefore we use this method as a baseline in our experiments.
Overall, among the related work, we argue that existing methods exhibit a number of Limitations (L1-L5). In summary, the model-dependent methods require training full embeddings first, limiting the maximum number of items that can be considered in the catalogue of the recommender system. On the other hand, methods which do not require training full embeddings first, such as Quotient Remainder, rely on simple heuristics and may result in unrelated items having similar representations. On the surface, the nearest related work to ours is VQ-Rec (Hou et al., 2023) as it also applies JPQ-style technique to sequential recommendation; however, similarly to JPQ, it relies on the availability of textual features and pre-trained language models. In contrast, our work’s main novelty is adapting JPQ to the scenario when (e.g., textual) side information is unavailable. In the next section, we describe JPQ (Zhan et al., 2021a), a method of embedding compression for information retrieval, which directly learns embeddings in compressed form, reducing GPU memory requirements. Then, in Section 4, we propose RecJPQ - an adaptation of JPQ to the sequential recommendation task, which successfully addresses Limitations L1-L5.
3. Product Quantisation and JPQ
Figure 2. Reconstruction of item embeddings in RecJPQ: Codebook length $m=3$, item embedding length $d=12$, number of centroids per split $b=256$.
We now describe Product Quantisation (PQ) and Joint Product Quantisation (JPQ), two methods which serve as a backbone for our method. Section 3.1 covers PQ, a classic embedding compression technique. Section 3.2 describes JPQ - a recently proposed information retrieval method that learns compressed embeddings directly instead of compressing them after the model training.
3.1. Product Quantisation
@@
Product quantisation (Gray, 1984; Jégou et al., 2011) is a well-cited method of compressing vectors used by many libraries, such as FAISS (Johnson et al., 2021) & nanopq [^9] https://github.com/matsui528/nanopq. Its main idea is to split a collection of $d$-dimensional vectors, $V$, into $m$ collections $V_{i};i=1..m$ of smaller vectors of $\frac{d}{m}$ dimensions each. The original vectors can be recovered back via concatenation: $V=concat(V_{1},V_{2},...V_{m})$. Product quantisation then clusters each $V_{i}$ into $b$ clusters (e.g. using the k-means algorithm (MacQueen, 1967)) and replaces each vector $v_{ij}$ with the centroid of the assigned cluster $c_{ij}\approx v_{ij}$. With this replacement, the original vectors collection can be approximated as a concatenation of the centroid matrices $C_{1}...C_{m}$, which are constructed from $V_{1}...V_{m}$ by replacing each vector $v_{ij}$ by the closest centroid $c_{ij}$:
(1) $V\approx concat(C_{1},C_{2},...C_{m})$
Note that in each centroid matrix $C_{i}$, there are, at most, $b$ different rows, as each row corresponds to one of the centroids of the clusters, so instead of storing full matrix $C_{i}$, we can store these unique rows in the separate tensor $\hat{C}_{i}$, whose elements $\hat{c}_{ij},j\in\{1..b\}$ correspond to the unique centroids. To compress the approximate embedding matrix defined by Equation (1), we need only store the ids of centroids. Overall, there are $m$ centroid sets $C_{i}$, so each vector can be encoded using $m$ integer codes, and each code can have $b$ different values; therefore, overall, this scheme can encode up to $b^{m}$ different vectors. The vector of centroid ids $g_{i}=\{g_{i1},g_{i2}...,g_{im}\}$ associated with the vector $v_{i}\in V$ is known as the code of the vector $v_{i}$ (Jégou et al., 2011). The number of centroids associated with an item $m$ is the length of the code. The table $G$ of codes associated with each vector from $V$ is also known as a codebook (Jégou et al., 2011).
Figure 2 illustrates the vector reconstruction process applied to item embeddings. For each centroid id $g_{ij}$ in the codes vector $g_{i}$ of an item $i$, we extract a centroid associated with this centroid id and then concatenate the centroids to obtain reconstructed item embeddings.
101010
The number of splits $m$ is usually considerably smaller than the original vector dimensions $d$: $m\ll d$ to achieve compression. The number of centroids per split, $b$, is typically a power $k$ of 256, so that the codes can be stored as $k$-byte integers. In this paper, for simplicity and following JPQ (Zhan et al., 2021a), we fix $k=1$, so each centroid id can be represented with a single byte; therefore, we only store $m$ bytes for each item. Even with fixed $k$ (and therefore $b$), we can adjust model capacity by controlling the number of centroids associated to each item, $m$, and the dimension of the centroid embeddings, $\frac{d}{m}$, by adjusting $d$. Figure 2 illustrates RecJPQ for $m=3$, $d=12$, $b=256$. Further, to illustrate the achieved compression, if embeddings are stored as 256-dimensional float32 vectors, a full-item embedding requires 1KB of memory. After compression using product quantisation with $m=8$ splits, we only need to store 8 bytes per item (0.78% of the original memory requirement). While some memory is also required to store the centroids themselves, this is a negligible for large datasets compared to the original vector requirements (see also Table 2 for an analysis of centroids memory requirements).

### 18 Dec 2023 v2

@@ Abstract.
Sequential Recommendation is a popular recommendation task that uses the order of user-item interaction to model evolving users’ interests and sequential patterns in their behaviour. Current state-of-the-art Transformer-based models for sequential recommendation, such as BERT4Rec and SASRec, generate sequence embeddings and compute scores for catalogue items, but the increasing catalogue size makes training these models costly. The Joint Product Quantisation (JPQ) method, originally proposed for passage retrieval, markedly reduces the size of the retrieval index with minimal effect on model effectiveness, by replacing passage embeddings with a limited number of shared sub-embeddings. This paper introduces RecJPQ, a novel adaptation of JPQ for sequential recommendations, which takes the place of item embeddings tensor and replaces item embeddings with a concatenation of a limited number of shared sub-embeddings and, therefore, limits the number of learnable model parameters. The main idea of RecJPQ is to split items into sub-item entities before training the main recommendation model, which is inspired by splitting words into tokens and training tokenisers in language models. We apply RecJPQ to SASRec, BERT4Rec, and GRU4rec models on three large-scale sequential datasets. Our results showed that RecJPQ could notably reduce the model size (e.g., 48$\times$ reduction for the Gowalla dataset with no effectiveness degradation). RecJPQ can also improve model performance through a regularisation effect (e.g. +0.96% NDCG@10 improvement on the Booking.com dataset). Overall, RecJPQ allows the training of state-of-the-art transformer recommenders in industrial applications, where datasets with millions of items are common.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 17th ACM International Conference on Web Search and Data Mining; March 4–8, 2024; Merida, Mexicobooktitle: Proceedings of the 17th ACM International Conference on Web Search and Data Mining (WSDM ’24), March 4–8, 2024, Merida, Mexicoprice: 15.00doi: 10.1145/3616855.3635821isbn: 979-8-4007-0371-3/24/03
@@ 1. Introduction
When a recommender has many items in the catalogue, various challenges arise in training the neural recommendation model. Firstly, the item embedding tensor may contain more model parameters than the rest of the model. For example, there are more than 800 million videos on YouTube (you, 2023). If a recommender model uses 128-dimensional embeddings, the whole item embeddings tensor will have more than 100 billion parameters, which is comparable with the number of parameters of the largest available machine learning models (Brown et al., 2020), even without accounting for the parameters of the model’s intermediate layers. This is a problem specific to recommender systems: in the related area of dense passage retrieval (Karpukhin et al., 2020; Khattab and Zaharia, 2020), passage embeddings are obtained by encoding passage text using a pre-trained language model; however, item side information, such as text, is not necessarily available in a typical recommender systems scenario; therefore, item embeddings should be directly learned from the interactions. Secondly, a large number of such trainable parameters also makes the model prone to overfitting. A third challenge caused by the large catalogue is the size of the output scores tensor (rightmost tensor in Figure 1): for example, in BERT4Rec, it contains a score for each item for every position, for every sequence in the training batch, so training BERT4Rec with more than 1 million items in the catalogue may be prohibitively expensive (Petrov and Macdonald, 2022a). This problem is typically solved using negative sampling, whereby instead of computing the full output tensor, the model computes scores for a small proportion of negative items not interacted by the user (e.g. SASRec (Kang and McAuley, 2018) uses one negative per positive). However, negative sampling comes with its own challenges (for example, it usually requires informative negative mining (Rendle, 2022)). Nevertheless, negative sampling is an orthogonal research direction, and in this paper, we use SASRec in cases where negative sampling is necessary. We refer the reader to our recent publication (Petrov and Macdonald, 2023), where we analysed the negative sampling problem in details. To summarise the challenges, a large embedding tensor increases the model size, slows model training down, and can necessitate further modelling tricks such as negative sampling, which bring their own challenges.
@@
There are some existing methods (Xia et al., 2023; Wang et al., 2022; Kang et al., 2020) for item embedding compression (we discuss these methods in Section 2). However, most of these methods compress the embedding tensor after the model is fully trained (including training the full embedding tensor). However, as argued above, training may be prohibitively expensive in large-catalogue recommender systems. Hence, this paper addresses the problem of a large item embedding tensor in sequential recommendation models at the training stage.
@@
To mitigate this problem, we propose a novel RecJPQ technique inspired by the success of a recent Joint Product Quantisation (JPQ) work (Zhan et al., 2021a) for passage retrieval. JPQ itself is based on Product Quantisation (PQ) (Jégou et al., 2011), a popular method of compressing vectors by splitting them into sub-embeddings and encoding them using a discrete item [^3] For consistency, we explain prior work using item ids instead of passage ids. codebook (the codebook maps from item ids to the associated sub-item ids; see Section 3.1 for the details). The main innovation of JPQ compared to the standard PQ method is that it learns the sub-item embeddings as part of the overall model training process. In contrast, PQ requires training the model first and only then compressing the embeddings (frequently, this second step uses external tools, such as FAISS (Johnson et al., 2021)). This means that JPQ does not need to keep the embedding matrix in memory during model training. We argue that this innovation is valuable for recommender systems. Indeed, as mentioned above, real-life recommender systems can have hundreds of millions of items in their catalogues and keeping full embeddings tensor in memory may be prohibitively expensive. This is particularly important for deep-learning-based sequential recommender systems because these models require keeping the whole model in GPU (or TPU) memory during training. GPU memory is costly even when compared to regular computer RAM.
@@
Unfortunately, it is hard to adapt JPQ to the recommendation scenario, as it is specific to textual information retrieval. In particular, JPQ assumes the existence of a pre-trained (passage retrieval) model and index, which it uses to assign items to sub-item ids (see more details in Section 3.2). These pre-trained models rarely exist in item recommendations. Hence, in RecJPQ, we experiment with performing the initial assignment of sub-ids using three different strategies: (i) discrete truncated SVD (sub-ids obtained by discretising the item representations obtained by an SVD decomposition of the user-item matrix), (ii) discrete BPR (sub-ids obtained by discretising the item embeddings obtained from BPR) and (iii) random assignments. We describe these assignment strategies in detail in Section 4.
RecJPQ is a model component that replaces traditional item embeddings in sequential recommender systems. In general, it can be applied to any recommender system based on item embeddings, but in this paper, we focus specifically on sequential models, as in these models, item embeddings comprise the biggest part of the model (e.g. sequential models usually do not have user embeddings). In contrast with existing methods, RecJPQ does not require training full uncompressed embedding and does not modify the original model loss function. Our experimentation on three datasets (see Section 5) demonstrates that RecJPQ can be successfully applied to different models, including SASRec (Kang and McAuley, 2018), BERT4Rec (Sun et al., 2019) and GRU (Hidasi et al., 2016; Petrov and Macdonald, 2022a), achieving a large factor of embeddings compression (e.g. 47.94$\times$ compression of SASRec on Gowalla) without any effectiveness degradation. Moreover, on 2 out of our 3 experimental datasets, applying RecJPQ increases model performance (e.g. +0.96% NDCG@10 on Booking.com dataset, significant improvement); we attribute these improvements to model regularisation.
In short, the contributions of this paper are as follows: (1) we propose RecJPQ, a novel technique for reducing the size of sequential recommendation models during training based on Joint Product Quantisation; (2) we propose three strategies for assigning sub-item ids to items, two of which (discrete truncated SVD and discrete BPR) assign similar codes to similar items, and one assigns codes randomly; (2) we perform an extensive experimental evaluation of RecJPQ on three datasets and show that RecJPQ allows reducing the models’ size without hindering the model performance. The rest of the paper is organised as follows: Section 2 introduces related work on embeddings compression and identifies limitations of existing methods; Section 3 covers Product Quantisation (PQ) and Joint Product Quantisation (JPQ) – the methods, which serve as the basis for our work; Section 4 introduces RecJPQ and covers sub-id assignment strategies for RecJPQ; in Section 5 we experimentally evaluate RecJPQ; Section 6 contains final remarks.
555
@@ 2. Related work
Table 1. Existing embedding compression methods. The desired method characteristics are highlighted in bold.
@@
This section covers existing work on compressing and discretising embeddings in recommender systems, identifies the limitations in existing work and positions our contributions in the context of existing methods. Table 1 summarises existing methods and positions RecJPQ, our proposed compression technique. The table highlights with boldface the desirable characteristics necessary for training a large-scale [^6] For simplicity, we say that a catalogue is ”large-scale” if it has more than 1 million items, as it becomes challenging to train recommender systems on that scale (Petrov and Macdonald, 2022a). sequential model, specifically: the method can be applied to work with different backbone sequential models, and does not require training full embeddings (as we work with the assumption that the full embeddings tensor does not fit into GPU memory); and we want the model to focus on item embeddings rather than embeddings of other entities, such as users or features. As illustrated in the table, the methods for compressing the models can be broadly divided into two groups: model dependent and model agnostic.
In the model-dependent methods (Lian et al., 2020; Xia et al., 2023), the embedding compression mechanism is integrated as a component into the recommendation model itself. Hence, the training architecture of these methods has to be aware of the compression, and the loss function includes components responsible for the embedding compression. For example, LightRec (Lian et al., 2020) uses the Deep Semantic Similarity Model (DSSM) (Huang et al., 2013) as the backbone model and uses an additional knowledge distillation component in the loss, which allows for learning compressed representations of the embeddings. However, DSSM is not a sequential model, and it is unclear whether or not the method can be adapted to the sequential recommendation task. Similarly, EODRec (Xia et al., 2023), uses SASRec (Kang and McAuley, 2018) as its backbone, one of the most popular sequential models based on the Transformer architecture (Vaswani et al., 2017). The loss function of EODRec also consists of four components, some of which are responsible for recommendation, and others are responsible for embedding compression. While EODRec’s backbone (SASRec) is an efficient model, in many cases, other models such as BERT4Rec exhibit better results (Sun et al., 2019; Petrov and Macdonald, 2022b). In general, while some model-dependent methods may reduce the size of a trained model, these methods have several limitations, which make them unsuitable for training sequential recommendation models with large catalogues. In particular:
@@
777
L2 Model-dependent methods usually require training (uncompressed) item embeddings and then use knowledge distillation or teacher-student techniques to obtain compressed representations of the embeddings. This approach substantially reduces the final model size, thereby helping inference on smaller devices, but requires a large amount of GPU memory while training, thereby limiting the overall number of items in the catalogue. For this reason, the main positioning for EODRec model (Xia et al., 2023) is the on-device recommendation: while the final model produced by this method is small, it requires storing full item embeddings while training. Post-training quantisation methods (Yao et al., 2022), which recently became popular to reduce the size of large language models via quantising their weights into lower-precision numbers (e.g. float16, or int8) also have this limitation – they need to have access to the full model before quantising. Similarly, Mixed Precision Training (Micikevicius et al., 2018) builds a smaller precision model, but it requires keeping full precision weights in memory. Placing the embeddings tensor into Approximate Nearest Neighbours (Johnson et al., 2021) or Hierarchical Navigable Small Worlds (Malkov and Yashunin, 2020) indexes also requires access to the full embedding tensor at model training time, and therefore also exhibits this limitation.
@@
On the other hand, the existing model-agnostic methods (Shi et al., 2020a, b) do not depend on the specific model architecture, and likewise do not add extra components to the loss functions. Typically, these methods implement a mechanism that takes the place of the embeddings tensor in the backbone model, and hence can be used with many models. However, on inspection of the relevant work, we elicit additional limitations of these methods:
@@
L4 Many methods are not designed for compressing item embeddings. For example, PreHash (Shi et al., 2020a) is a method specific for compressing user embeddings (i.e. it uses the user’s history to construct user embeddings). The method uses an attention network over the history of user interactions. Adapting this network structure for items embeddings is a hard task: a user may only interact with a few items; in contrast, a popular item may be interact with by millions of users. The attention mechanism depends quadratically on the sequence length, and therefore, applying it to users who interacted with a popular item would be prohibitively expensive.
@@
L5 Finally, many methods lack structure in compressed item representations. This leads to situations where unrelated items have similar representations and, conversely, when similar items have dissimilar representations. Both these cases may limit the models’ generalisation ability and hinder the models’ performance. One example of such unstructured methods is hashing-based Quotient Remainder (Shi et al., 2020b), which compresses embeddings of categorical features (e.g., genres). Quotient Remainder assigns feature codes based on the quotient and the remainder of the division of the feature id by some number. When applied to item ids (items can also be seen as categorical features), the quotient and the remainder are unrelated to the item characteristics. Hence, similar items are unlikely to have similar codes. Nevertheless, Quotient Remainder is one of the few methods that can be used to train a model on a large-scale dataset, and therefore, we use this method as a baseline in our experiments.
@@
Overall, among the related work, we argue that existing methods exhibit a number of Limitations (L1-L5). In summary, the model-dependent methods require training full embeddings first, limiting the maximum number of items that can be considered in the catalogue of the recommender system. On the other hand, methods which do not require training full embeddings first, such as Quotient Remainder, rely on heuristics and may result in unrelated items having similar representations. On the surface, the nearest related work to ours is VQ-Rec (Hou et al., 2023) as it also applies JPQ-style technique to sequential recommendation; however, similarly to JPQ, it relies on the availability of textual features and pre-trained language models. In contrast, our work’s main novelty is adapting JPQ to the scenario when (e.g., textual) side information is not available. In the next section, we describe Product Quantisation – a vector compression method and JPQ (Zhan et al., 2021a), a method of embedding compression for information retrieval, which directly learns embeddings in compressed form, reducing GPU memory requirements. Then, in Section 4, we propose RecJPQ – an adaptation of JPQ to the sequential recommendation task, which successfully addresses Limitations L1-L5.
@@
3. Product Quantisation and JPQ
We now describe Product Quantisation (PQ) and Joint Product Quantisation (JPQ), two methods that serve as backbones for our method. Section 3.1 covers PQ, a classic embedding compression technique. Section 3.2 describes JPQ – a recently proposed information retrieval method that learns compressed embeddings directly instead of compressing them after the model training.
3.1. Product Quantisation
Product quantisation (Gray, 1984; Jégou et al., 2011) is a well-cited method of compressing vectors used by many libraries, such as FAISS (Johnson et al., 2021) & nanopq (Yusuke Matsui, 2023). Its main idea is to split a collection of $d$-dimensional vectors, $V$, into $m$ collections $V_{i};i=1..m$ of smaller vectors of $\frac{d}{m}$ dimensions each. The original vectors can be recovered back via concatenation: $V=concat(V_{1},V_{2},...V_{m})$. Product quantisation then clusters each $V_{i}$ into $b$ clusters (e.g. using the k-means algorithm (MacQueen, 1967)) and replaces each vector $v_{ij}$ with the centroid of the assigned cluster $c_{ij}\approx v_{ij}$. With this replacement, the original vectors collection can be approximated as a concatenation of the sub-vector matrices $C_{1}...C_{m}$, which are constructed from $V_{1}...V_{m}$ by replacing each vector $v_{ij}$ by the closest cluster centroid $c_{ij}$: $V\approx concat(C_{1},C_{2},...C_{m})$. In each sub-vector matrix $C_{i}$, there are, at most, $b$ different rows, as each row corresponds to one of the centroids of the clusters, so instead of storing full matrix $C_{i}$, we can store these unique rows in the separate tensor $\hat{C}_{i}$, whose elements $\hat{c}_{ij},j\in\{1..b\}$ correspond to the unique sub-vectors. To compress the approximate embedding matrix, we need only store the ids of the sub-vectors (sub-ids) for each item. Overall, there are $m$ sub-vector sets $C_{i}$, so each vector can be encoded using $m$ integer codes, and each code can have $b$ different values; therefore, overall, this scheme can encode up to $b^{m}$ different vectors. The vector of sub-item ids $g_{i}=\{g_{i1},g_{i2}...,g_{im}\}$ associated with the vector $v_{i}\in V$ is known as the code of the vector $v_{i}$ (Jégou et al., 2011). The number $m$ of sub-item ids associated with each item is called the length of the code. The table $G$ of codes associated with each vector from $V$ is also known as a codebook (Jégou et al., 2011).
Figure 2 illustrates the vector reconstruction process applied to item embeddings. For each sub-item id $g_{ij}$ in the codes vector $g_{i}$ of an item $i$, we extract a sub-item embedding associated with this sub-item id and then concatenate the sub-item embeddings to obtain reconstructed item embeddings.
Figure 2. Reconstruction of item embeddings using Product Quantisation: Codebook length $m=3$, item embedding length $d=12$, number of sub-ids per split $b=256$.
999
The number of splits $m$ is usually considerably smaller than the original vector dimensions $d$: $m\ll d$ to achieve compression. The number of sub-embeddings per split, $b$, is typically a power $k$ of 256, so that the codes can be stored as $k$-byte integers. In this paper, for simplicity and following JPQ (Zhan et al., 2021a), we fix $k=1$, so each sub-item id can be represented with a single byte; therefore, we only store $m$ bytes for each item. Even with fixed $k$ (and therefore $b$), we can adjust model capacity by controlling the number of sub-item ids associated with each item, $m$, and the dimension of the sub-item embeddings, $\frac{d}{m}$, by adjusting $d$. Figure 2 illustrates PQ applied to item embeddings for $m=3$, $d=12$, $b=256$. Further, to illustrate the achieved compression, if embeddings are stored as 256-dimensional float32 vectors, a full-item embedding requires 1KB of memory. After compression using product quantisation with $m=8$ splits, we only need to store 8 bytes per item (0.78% of the original memory requirement). While some memory is also required to store the sub-item embeddings themselves, this is negligible for large datasets compared to the original vector requirements (see also Table 2 for an analysis of sub-item embeddings memory requirements).